What is the Kirchhoff's junction rule equation for
Answer #1

Q1.a )

kirchoff's current law states that:

total current entering at a node=total current leaving a node

here total current entering=0 A

total current leaving =I1+I2+I3

hence KCL states that:

I1+I2+I3=0

hence option d is correct.


Q1.b)

time constant of the circuit=R*C=2*2*10^(-6)=4 us

hence several seconds is enough for the circuit to reach steady state.

at steady state, capacitor acts like an open circuit

hence after several seconds of switch being closed, capacitor will be open circuit

and hence current in the circuit=0

capacitor will be charged to the connected potential=10 volts

hence ammeter reading is 0 A and voltmeter reading is 10 volts.

hence option a is correct.

Q1.c)

as inductor behaves like an open circuit when the switch is just closed,

initial current=0 A

after steady state is reached, inductor behaves like a short circuit

hence current in the circuit=V/R

the current will rise exponentially

hence current at any time t is given as

i(t)=(V/R)*(1-exp(-t/T))

where T=L/R=time constant

graph:

current vs time 2.5 1.5 05- 0.005 0.01 0.015 0.02 0.04 0.025 time (in seconds 0.03 035 0.045 0.05

Q1.d)maximum current=steady state current=V/R=24/10=2.4 A